# Two Channel Multi impurity Kondo model: RKKY induced criticality

### v4

@@ I.4 Multichannel Kondo lattice
[…]
$\Delta_{i,j}$ is a matrix in cluster site indices i j, and is diagonal in channel indices. It is determined by the cluster DMFT self consistency condition. To demonstrate the relevance of the new fixed point, we use the Dynamical Cluster Approximation (DCA) Hettler et al. (1998) which makes the hybridization matrix cyclical, and its Fourier transform diagonal in the cluster momenta $K_{i}$. This makes it straightforward to identify the channels, which are now labeled by the cluster momenta. The DCA can also be understood in terms of a coarse graining of the lattice Brillouin zone into N patches of the same size, with the cluster momenta $K_{i}$ labeling each patch.
@@
Under the assumption that the impurity self energy $\Sigma_{imp}$ of the problem under consideration is a not divergent at low frequency, analytic continuation of Eqn. 42 gives a finite value $-Im\Delta(\omega+i\delta)$ at zero frequency, coming from the k integration in the patches which cross the fermi level. This statement is valid as long as the conduction bands are sufficiently broad so that the real part of the self energy $\Sigma_{imp}$ does not shift zeros of the denominator in the sum over k in Eqn. 42 beyond the edge of the band of the conduction electrons. Having shown that that $-Im\Delta(\omega+i\delta)\neq 0$ we now have the spin sector impurity model considered in the text, together with a decoupled charge and channel sector, and the spin spin correlation functions of the cluster DMFT solution at low energies are given by the results presented in the main text. The evaluation of correlation functions involving the charge sector are left for future studies.
Acknowledgments: We would like to thank Piers Coleman for fruitful discussions. This work was supported by Office of Basic Energy Sciences, Material Sciences and Engineering Division, U.S. Department of Energy (DOE) under Contracts No. DE-SC0012704 and by the National Science Foundation (GK) under DMR-1733071.

### v5

@@ I.3 Multichannel Kondo lattice
$S_{\text{conduction}}=\sum_{m=1}^{2}\sum_{\sigma=1}^{2}\sum_{i=1}^{N}\sum_{j=1}^{N}\int d\tau\int d\tau^{\prime}\,\psi_{mj\sigma}^{\dagger}(\partial_{\tau}+\Delta_{i,j}(\tau-\tau^{\prime}))\psi_{mj\sigma},$ (39)
$S_{\text{impurity}}=\sum_{\sigma=1}^{2}\int d\tau\,f_{\sigma}^{\dagger}(\partial_{\tau}+\lambda)f_{\sigma}-\lambda,$ (40)
$S_{\text{interaction}}=J\sum_{m=1}^{2}\sum_{j=1}^{N}\int d\tau\,\mathbf{S}\cdot\mathbf{s}_{jm}(0)$ (41)
$\Delta_{i,j}$ is a matrix in cluster site indices i j, and is diagonal in channel indices. It is determined by the cluster DMFT self consistency condition. To demonstrate the relevance of the new fixed point, we use the Dynamical Cluster Approximation (DCA) [45] which makes the hybridization matrix cyclical, and its Fourier transform diagonal in the cluster momenta $K_{i}$. This makes it straightforward to identify the channels, which are now labeled by the cluster momenta. The DCA can also be understood in terms of a coarse graining of the lattice Brillouin zone into N patches of the same size, with the cluster momenta $K_{i}$ labeling each patch.
@@
Under the assumption that the impurity self energy $\Sigma_{imp}$ of the problem under consideration is a not divergent at low frequency, analytic continuation of Eqn. 42 gives a finite value $Im\Delta(\omega+i\delta)$ at zero frequency, coming from the k integration in the patches which cross the fermi level. This statement is valid as long as the conduction bands are sufficiently broad so that the real part of the self energy $\Sigma_{imp}$ does not shift zeros of the denominator in the sum over k in Eqn. 42 beyond the edge of the band of the conduction electrons. Having shown that that $Im\Delta(\omega+i\delta)\neq 0$ we now have the spin sector impurity model considered in the text, together with a decoupled charge and channel sector, and the spin spin correlation functions of the cluster DMFT solution at low energies are given by the results presented in the main text. The evaluation of correlation functions involving the charge sector are left for future studies.
Acknowledgments: We would like to thank Piers Coleman and Andriy Nevidomskyy for fruitful discussions. This work was supported by Office of Basic Energy Sciences, Material Sciences and Engineering Division, U.S. Department of Energy (DOE) under Contracts No. DE-SC0012704 and by the National Science Foundation (GK) under DMR-1733071.